OpenStudy (anonymous):

[calculus] differentiate: f(x) = (lnx) / x^3

OpenStudy (anonymous):

quotient rule for this one \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\ln(x), f'(x)=\frac{1}{x}, g(x)=x^3,g'(x)=3x^2\] plug and chug

OpenStudy (kainui):

\[f(x)=x^{-3}lnx\] Or use exponent rules to make this a nice product rule question, I always forget which is negative in the quotient rule.

OpenStudy (anonymous):

Thanks guys!